OpenStudy (anonymous):

SUMMER MATH PACKET!! Please help check my answers?? Thanks! Directions: Solve. 1. 3y-9<6+6y my answer: y<5 2. 1/4 (x-2) - 1/5 (x-3) = 1/2 my answer: x=8 3. -3x+3y=-18 4x-y=5 my answer: x=-13/9 and y=-67/9 4. 2x-3y=8 5x+2y=1 my answer: x=1 and y=-2 5. 2+x/5=5 my answer: x=23 6. (x+3)(x-9)=0 my answer: (-3,9)

OpenStudy (anonymous):

nvm I have the first two corrected but still need 3-6

OpenStudy (anonymous):

k

OpenStudy (anonymous):

3rd one is x=-1/3 and y=-17/3

OpenStudy (anonymous):

how did you get there?

OpenStudy (anonymous):

4th is right

OpenStudy (anonymous):

-3x+3y=-18 -x+y=-6.....eqn 1 simply add this to eqn2 4x-y=5 we get 3x=-1 or x=-1/3 put this back to eqn1 and i get y=-17/3

OpenStudy (anonymous):

i have doubt in 5th is it 2+(x/5) or (2+x)/5 ?

OpenStudy (anonymous):

(2+x)/5

OpenStudy (anonymous):

y= -19/3

OpenStudy (anonymous):

is this answer @greeneyes<3 ?

OpenStudy (anonymous):

@heybhai yes it is y=-19/3 not -17/3 thank you for your help! :D

OpenStudy (anonymous):

welcome and I thing I wanna admit u r very beautiful.